On 7/1/20, Cenwein Corp. recorded the purchase of manufacturing equipment as follows:

Equipment Expense                                                                $850,000

            Cash                                                                                                    $850,000

Cenwein anticipates using the equipment for 8 years and a residual value of $26,000 at the end of that time.

  1. If the company does not make any correcting or adjusting entries during 2020 or 2021, determine the direction and magnitude of the error under US GAAP accrual accounting standards for assets, liabilities, and net income and equity.

Asse   Assets 12/31/20

Liabilities 12/31/20

Net Income 2020

Equity 12/31/20

Asset Assets 12/31/21

Liabilities 12/31/21

Net Income 2021

Equity 12/31/21
Answer #1

Answer:

Assets 12/31/20 Liabilities 12/31/20 Net Income 2020 Equity 31/12/20
$7,47,000 0 -$1,03,000 0
Assets 12/31/21 Liabilities 12/31/21 Net income 2021 Equity 12/31/21
$6,44,000 0 -$1,03,000 0

Assets 12/31/20 = $850000-$103000 =$7,47,000/-

Assets 12/31/21 =$747000-103000 =$6,44,000/-

Dep. = 850000-26000/8

= $103,000/-